# Ultra-fast Digital DPC Yielding High Spatio-Temporal Resolution for Low-Dose Phase Characterisation

### v1

@@ Abstract
In the scanning transmission electron microscope, both phase imaging of beam-sensitive materials and characterisation of a material’s functional properties using in-situ experiments are becoming more widely available. As the practicable scan speed of 4D-STEM detectors improves, so too does the temporal resolution achievable for both differential phase contrast (DPC) and ptychography. However, the read-out burden of pixelated detectors, and the size of the gigabyte to terabyte sized data sets, remain a challenge for both temporal resolution and their practical adoption. In this work, we show that a high-fidelity DPC phase reconstruction can be achieved from both annular segmented detectors or pixelated arrays with relatively few elements using signal digitisation. Unlike conventional analog data, even at the fastest scan speeds, phase reconstructions from digitised DPC-segment images yield reliable data. Finally, dose fractionation by fast scanning and multi-framing allows for post-process binning of frame streams to balance signal-to-noise ratio and temporal resolution for low-dose phase imaging for in-situ experiments.
@@ Introduction
Scanning transmission electron microscopy (STEM) is a commonly used technique for the structural, functional and chemical characterisation of materials at very high spatial resolution. In recent decades, the traditional monolithic bright field (BF) and annular dark field (ADF) detectors have started to be superseded by other geometries of detectors enabling new imaging techniques. One of these is the development of high dynamic range cameras capable of capturing the entire two-dimensional (2D) convergent beam electron diffraction (CBED) pattern for every 2D scan point on the specimen. This results in a four-dimensional (4D) dataset, and the detectors are therefore aptly referred to as 4D-STEM or pixelated detectors. These 4D cameras are becoming increasingly popular for many different types of materials characterisation with imaging methods that require position-sensitive detectors to elucidate the internal structure of the bright field disk and overlapping regions with scattered disks.
@@
Figure 1: Illustration of the practicable scan speed (in $\frac{1}{\delta_{t}}$) versus number of segments/pixels for various commercially available or researched detectors. Several detectors may be operated in different binning modes, where higher binning allows for faster data readout due to lower pixel numbers. All the data has been acquired from data-sheets or publications reporting the highest practicable readout speed from the detectors [bardelloni_new_2000, ponchut_maxipix_2011, ballabriga_medipix3_2011, ishikawa_high_2020, ballabriga_introduction_2020, wang_electron_2021, philipp_very-high_2022, ercius_4d_2023, grieb_gan_2024, ishikawa_real-time_2024]. Where several bit-depth modes are available, the main reported readout speed is quoted. Data used may be found at [detector-frame-rate-data].
@@
Recent advances in ptychographic reconstruction reaching information limits of less than $0.5$ uses both a high dwell time and high beam current which is challenging for samples that damage under beam irradiation [nguyen_achieving_2024]. The core assumption for ptychographic reconstruction that the sample is static poses challenges if the goal is to image dynamic processes in-situ. Especially when the minimum practicable dwell time for pixelated detectors is still typically found in the $10-1000$ µs regime. The assumption of the static object phase would only hold for samples that vary on time scales much higher than a few seconds, depending on the number of scan points (field of view) and the resolution required.
@@
Since pixelated detectors were first used to acquire images of the CBED in a STEM scan in the late 90’s with the first Medipix-detector, the data rate has been increasing with the development of faster and higher pixel-number detectors; these include the Quantum Detectors MerlinEM (Medipix3), TFS EMPAD (2015), Dectris Arina (2020), and the DirectElectron CeleritasXS (2023). To contextualise the speed and pixelation of various detectors, these parameters are plotted in Figure 1. Most pixelated detectors support hardware binning of pixels to get a higher read-out speed at the expense of a lower CBED sampling. This binning leads to a lower read-out delay and data set size, and facilitates higher real-space scan speeds and for this reason some detectors in Figure 1 are represented by multiple points. At the time of writing, in terms of frame-rate for fully pixelated 4D cameras, the state of the art is the Dectris Arina detector with a minimum real-space dwell time of $10$ µs. [stroppa_stem_2023].
@@
While 4D-STEM enables high-resolution imaging in both the image and detector plane, there are also position-sensitive detectors with a lower number of discrete segments which have been shown to yield high-quality image data for phase contrast techniques such as ptychography, optimum bright field (OBF), and differential phase contrast (DPC). Specifically, DPC has long been established as a powerful tool for probing both magnetic and electric fields inside samples, including the local electric fields surrounding individual atoms [rose_phase-contrast_1974, dekkers_differential_1974, lohr_differential_2012, shibata_differential_2012]. There are several examples where the functional properties of specimens, such as ferroelectric domain formation or internal magnetic fields, have been characterised with fewer numbered segmented detectors [ooe_ultra-high_2021, shibata_differential_2012, chen_imaging_2020]. Furthermore, a significant amount of noteworthy high-quality DPC imaging has been done with as low as four annular segments [shibata_new_2010, shibata_differential_2012, close_towards_2015, shibata_electric_2017, ishikawa_spatial_2023, grieb_gan_2024]. Similarly, almost a decade ago, Yang et.al. showed that for single-side-band (SSB) ptychography the signal-to-noise ratio (SNR) of the reconstructed image does not increase significantly with increased detector plane pixelation beyond around 16x16 [yang_efficient_2015]. And just a few years ago, Ooe et.al. demonstrated impressive low-dose phase reconstruction with the OBF technique with a 16 segmented annular detector [ooe_ultra-high_2021].
@@
In this work, we present the first demonstration of DPC imaging utilising a fast hardware-digitised electron counting detector signal. We also show that digitisation allows for much faster usable frame-rates and paves the way for low-dose phase characterisation, with potential use cases in beam-sensitive and low atomic number materials, as well as in-situ experiments.
@@
Here, $ref$ is the infinite dose reference image and $expt$ is the finite dose image [yang_efficient_2015]. Figure 3 shows that the PSNR of the images in Figure 2 reaches a plateau once the detector pixelation is 16x16 pixels. This calculation was repeated for several different finite doses, calculating both iCOM and magnitude of COM signals, and they all show the same result of minimal improvement of the PSNR for higher number of pixels in the detector beyond 16x16 (see Supplementary Information). Since the improvement of iCOM phase retrieval is minimal for higher sampling of the CBED, the detector pixel count can be kept low to allow for faster scanning. This is not to say that larger format cameras are not useful, in fact they are used widely for strain mapping [wang_autodisk_2022], PACBED [lebeau_position_2010, hwang_nanoscale_2012], crystallography characterisation with scanning electron (precession) diffraction (S(P)ED) [rauch_automated_2010, midgley_precession_2015, jeong_automated_2021], and more exotic applications like multi-beam electron diffraction [hong_multibeam_2021] and patterned probe Bragg measurements [zeltmann_patterned_2020]. Rather, it is just that for their use in DPC and iCOM, the additional pixelation offers no advantage.
@@
4D-STEM detection affords the opportunity to detect all the electrons from $0-\alpha_{max}$, where $\alpha_{max}$ is the outer collection angle set by the camera length, often at high resolution. While this offers very precise estimates for the COM signal, there are several other works that suggest losing the center of the CBED, as we do when using an annular segmented detector, causes minor loss to the final iCOM signal [muller_atomic_2014, shibata_differential_2012, majert_high-resolution_2015, shibata_electric_2017, seifer_flexible_2021, kohl_optimized_2023]. A similar result has been shown for ptychography [song_hollow_2018]. It has previously been shown that the PSNR of single side-band (SSB) ptychographic reconstruction of an isolated carbon atom increases very modestly for higher detector pixelation than 16x16 pixels [yang_efficient_2015]. Even segmented annular detectors of 4 and 16 segments can provide comparable ptychographic reconstructions. This further demonstrates how increasing pixelation of the detector is not necessarily better for phase contrast imaging. We may therefore reap the benefits of fewer element detectors (be they segments or pixels), to enable fast scanning and multi-frame to image with high spatial and temporal resolution.
@@
Using a probe-corrected JEOL Grand ARM operated at 300kV, single crystals of [100] oriented SrTiO3 were imaged with a 5.2 pA beam current. Several fast dwell times in the range 50 ns to 1 µs were collected with both with analog and digital signal recording. The microscope used is equipped with fast low-inductance scan coils allowing dwell times as short as $50$ ns to be reached [ishikawa_high_2020]. For the DPC imaging we used a segmented annular all-field (SAAF) detector, which has 16 segments spread over four concentric rings with four segments each [shibata_new_2010]. Performing the in-hardware digitisation has to be done on a per-segment basis and thus one channel per segment must be provided. For the experiments presented here, a four channel Pulse digitiser produced by turboTEM was used and the second ring from the center was chosen for the DPC imaging (see Supplementary Information). Digitisation of 8, 12, or 16 channels is possible by using additional Pulse digitisers in parallel so long as the scan unit can record their outputs (in this case a Gatan Digiscan 3). The outer collection angle of the detector annulus used was $32$ mrad and the inner collection angle $16$ mrad, for an $\alpha=30$ mrad beam.
Figure 4: Single pulse from the SAAF detector with the 10-90% rise time and 90-10% decay time indicated. The vertical, dotted lines are the pixel boundaries for a $250$ ns dwell time, and the stepped gray line indicates the normalised pixel intensity for each pixel with this dwell time. The resulting image pixels are shown in the diagram below, for dwell times $\delta_{t}=250,100,50$ ns respectively.
@@
The in-hardware digitisation calculates the gradient of the PMT signal and thresholds this to count the electron events. This results in a digitised signal which has a true zero noise floor, and where each electron is detected in the correct pixel only and all have the same intensity of unity. Digitising the signal from scintillator detectors has previously been demonstrated to significantly improve signal to noise ratio in HAADF images [mullarkey_development_2021, peters_electron_2023]. Since the digitised pulse appears in the correct temporal position in the PMT stream it also appears in the correct spatial position in the image, without affecting subsequent pixels. As a result, all streaking artefacts are removed. Incidentally, there is some unavoidable few nanoseconds progressing time required by the digitisation hardware, but this manifests as a uniform apparent translation of the sample to the right by some constant fraction of a pixel width. Since scintillator detector surfaces are typically inhomogenous, the electrons may be detected at different intensity depending on where on the detector surface they hit [macarthur_how_2014, sang_characterizing_2016]. Consequently, digitisation of the analog signal ensures a more homogeneous detector response, and yields an image that is quantitative, in number of electrons [macarthur_how_2014, mullarkey_how_2023]. This ensures that detector inhomogeneity does not affect the COM calculations, and that denoising algorithms developed for quantised signals may be used on the individual segment images before computing the COMs (although no denoising was used in this study). As a final advantage, having digitised signals makes quantitative comparison with simulations easier.
@@
Figure 6: PSNR in iCOM images for various summing of multi-frame stacks. Time per image is then given as the combined time it took to acquire all the frames, e.g. for 20 summed frames the time per image is $20\cdot T_{\mathrm{Frame}}$. The images on the right are the result of summing 20 frames for the three different dwell times $\delta_{t}=0.2,0.1,0.05$ µs, without and with a high-pass filter respectively.

### v2

@@ Abstract
In the scanning transmission electron microscope, both phase imaging of beam-sensitive materials and characterisation of a material’s functional properties using in-situ experiments are becoming more widely available. As the practicable scan speed of 4D-STEM detectors improves, so too does the temporal resolution achievable for both differential phase contrast (DPC) and ptychography. However, the read-out burden of pixelated detectors, and the size of the gigabyte to terabyte sized data sets, remain a challenge for both temporal resolution and their practical adoption. In this work, we combine ultra-fast scan coils and detector signal digitisation to show that a high-fidelity DPC phase reconstruction can be achieved from an annular segmented detector. Unlike conventional analog data phase reconstructions from digitised DPC-segment images yield reliable data, even at the fastest scan speeds. Finally, dose fractionation by fast scanning and multi-framing allows for post-process binning of frame streams to balance signal-to-noise ratio and temporal resolution for low-dose phase imaging for in-situ experiments.
@@ Introduction
Scanning transmission electron microscopy (STEM) is a commonly used technique for the structural, functional and chemical characterisation of materials at very high spatial resolution. In recent decades, the traditional monolithic bright field (BF) and annular dark field (ADF) detectors have started to be superseded by other geometries of detectors enabling new imaging techniques. One of these is the development of high dynamic range cameras capable of capturing the entire two-dimensional (2D) convergent beam electron diffraction (CBED) pattern for every 2D scan point on the specimen. This results in a four-dimensional (4D) dataset, and the detectors are therefore aptly referred to as 4D-STEM or pixelated detectors. These 4D detectors are becoming increasingly popular for many different types of materials characterisation with imaging methods that require position-sensitive detectors to elucidate the internal structure of the bright field disk and overlapping regions with scattered disks.
@@
Figure 1: Illustration of the practicable scan speed (in $\frac{1}{\delta_{t}}$) versus number of segments/pixels for various commercially available or researched detectors. Several detectors may be operated in different binning modes, where higher binning allows for faster data readout due to lower pixel numbers. All the data for the 4D detectors has been acquired from data-sheets or publications reporting the highest practicable readout speed from the detectors [bardelloni_new_2000, ponchut_maxipix_2011, ballabriga_medipix3_2011, tate_high_2016, ishikawa_high_2020, ballabriga_introduction_2020, wang_electron_2021, philipp_very-high_2022, ercius_4d_2023]. For the detectors marked with * the scan speed is taken from publications and does not necessarily represent the highest possible speed achievable [grieb_gan_2024, ishikawa_real-time_2024]. Where several bit depth modes are available, the main reported readout speed is quoted (for details see Supplementary Information). Data used may be found at [detector-frame-rate-data].
@@
Recent advances in ptychographic reconstruction reaching information limits of less than $0.5$ uses both a high dwell time and high beam current which is challenging for samples that damage under beam irradiation [nguyen_achieving_2024]. The core assumption for ptychographic reconstruction that the sample is static poses challenges if the goal is to image dynamic processes in-situ. Especially when the minimum practicable dwell time for pixelated detectors is still typically found in the $10-1000$ µs regime. The assumption of the static object phase would only hold for samples that vary on time scales much higher than a few seconds, depending on the number of scan points (field of view) and the resolution required. Still, it should be mentioned that with iterative ptychography and defocused probes the dose delivered to the sample may be reduced and temporal resolution improved due to the increased sample-plane step size [rodenburg_phase_2004, maiden_improved_2009, zhou_low-dose_2020].
@@
Since pixelated detectors were first used to acquire images of the CBED in a STEM scan in the late 90’s with the first Medipix-detector, the data rate has been increasing with the development of faster and higher pixel-number detectors; these include the TFS EMPAD (2015), Quantum Detectors MerlinEM (Medipix3, 2020), DirectElectron Celeritas (2021), and Dectris Arina (2022). To contextualise the speed and pixelation of various detectors, these parameters are plotted in Figure 1. Most pixelated detectors support hardware binning of pixels to get a higher read-out speed at the expense of a lower CBED sampling. This binning leads to a lower read-out delay and data set size, and facilitates higher real-space scan speeds and for this reason some detectors in Figure 1 are represented by multiple points. Further, the type and dynamic ranges of the commercially available 4D-STEM detectors in Figure 1 are summarised in Table S1 in the Supplementary Information. For several detectors the frame-rate may be increased by choosing a lower bit depth, essentially sacrificing dynamic range for speed. Still, bit depth does not tell the full story of the dynamic range achievable by the detector since some detectors encode more information than intensity recorded in the pixel in their output, with the result that a higher dynamic range is supported. At the time of writing, in terms of frame-rate for fully pixelated 4D detectors, the state of the art is the Dectris Arina detector with a minimum real-space dwell time of $10$ µs. [stroppa_stem_2023].
@@
While 4D-STEM enables high-resolution imaging in both the image and detector plane, there are also position-sensitive detectors with a lower number of discrete segments which have been shown to yield high-quality image data for phase contrast techniques such as ptychography, optimum bright field (OBF), and differential phase contrast (DPC). Specifically, DPC has long been established as a powerful tool for probing both magnetic and electric fields inside samples, including the local electric fields surrounding individual atoms [rose_phase-contrast_1974, dekkers_differential_1974, lohr_differential_2012, shibata_differential_2012]. Although reliable DPC signals depend on minimal dynamical scattering, and thus requires a very thin sample, there are several examples where the functional properties of specimens, such as ferroelectric domain formation or internal magnetic fields, have been characterised with fewer numbered segmented detectors [ooe_ultra-high_2021, shibata_differential_2012, chen_imaging_2020]. Furthermore, a significant amount of noteworthy high-quality DPC imaging has been done with as low as four annular segments [shibata_new_2010, shibata_differential_2012, close_towards_2015, shibata_electric_2017, ishikawa_spatial_2023, grieb_gan_2024]. Similarly, almost a decade ago, Yang et. al. showed that for single-side-band (SSB) ptychography the signal-to-noise ratio (SNR) of the reconstructed image does not increase significantly with increased detector plane pixelation beyond around 16x16 [yang_efficient_2015]. And just a few years ago, Ooe et. al. demonstrated impressive low-dose phase reconstruction with the OBF technique with a 16 segmented annular detector [ooe_ultra-high_2021].
@@
In this work, we combine for the first time ultra-fast scan coils [ishikawa_high_2020] and in-hardware digitisation of segmented scintillator-based detector [peters_electron_2023] to show that digitisation allows for much faster usable frame-rates and paves the way for low-dose phase characterisation, with potential use cases in beam-sensitive and low atomic number materials, as well as in-situ experiments.
@@
Here, $ref$ is the infinite dose reference image and $expt$ is the finite dose image [yang_efficient_2015]. Figure 3 shows that the PSNR of the images in Figure 2 reaches a plateau once the detector pixelation is 16x16 pixels. This calculation was repeated for several different finite doses, calculating both iCOM and magnitude of COM signals, and they all show the same result of minimal improvement of the PSNR for higher number of pixels in the detector beyond 16x16 (see Supplementary Information). Since the improvement of iCOM phase retrieval is minimal for higher sampling of the CBED, the detector pixel count can be kept low to allow for faster scanning. This is not to say that larger format cameras are not useful, in fact they are used widely for strain mapping [wang_autodisk_2022], PACBED [lebeau_position_2010, hwang_nanoscale_2012], crystallography characterisation with scanning electron (precession) diffraction (S(P)ED) [rauch_automated_2010, midgley_precession_2015, jeong_automated_2021], and more exotic applications like multi-beam electron diffraction [hong_multibeam_2021] and patterned probe Bragg measurements [zeltmann_patterned_2020]. Rather, it is just that for their use in DPC and iCOM, the additional pixelation offers no advantage. Still, regardless of whether a pixelated or segmented detector is used, DPC places strong limitations on the experimental parameters like sample thickness, bending and the microscope operator’s ability to align the beam to zone axis. Because dynamical scattering and zone axis changes can produce a DPC signal indistinguishable from the DPC signal expected from an electric or magnetic field, careful considerations must be given to both sample preparation and the setup of the experiment [maclaren_origin_2015, burger_influence_2020].
@@
4D-STEM detection affords the opportunity to detect all the electrons from $0-\alpha_{max}$, where $\alpha_{max}$ is the outer collection angle set by the camera length, often at high resolution. While this offers very precise estimates for the COM signal, there are several other works that suggest losing the center of the CBED, as we do when using an annular segmented detector, causes minor loss to the final iCOM signal [muller_atomic_2014, shibata_differential_2012, majert_high-resolution_2015, shibata_electric_2017, seifer_flexible_2021, kohl_optimized_2023]. A similar result has been shown for ptychography [song_hollow_2018]. It has previously been shown that the PSNR of single side-band (SSB) ptychographic reconstruction of an isolated carbon atom increases very modestly for higher detector pixelation than 16x16 pixels [yang_efficient_2015]. Even segmented annular detectors of 4 and 16 segments can provide comparable ptychographic reconstructions. This further demonstrates how increasing pixelation of the detector is not necessarily better for phase contrast imaging. We may therefore reap the benefits of fewer element detectors (be they segments or pixels), to enable fast scanning and multi-frame to image with high spatial and temporal resolution. It should however be mentioned that low-dose ptychography is achievable and routinely performed using iterative ptychographic reconstruction algorithms [rodenburg_phase_2004, maiden_improved_2009, song_atomic_2019, zhou_low-dose_2020, li_4d-stem_2022]. These experiments often use highly defocused probes allowing large sample step sizes, which also improves the achievable temporal resolution. Additionally, more advanced methods like multi-slice ptychography and tilt-corrected BF-STEM are becoming more and more popular and showing promising results [maiden_ptychographic_2012, varnavides_simultaneous_2023, ribet_uncovering_2024, oleary_three-dimensional_2024]. For such techniques, increased pixelation of the detector beyond 16x16 is advantageous.
@@
Figure 4: Single pulse from the SAAF detector with the 10-90% rise time and 90-10% decay time indicated. The vertical, dotted lines are the pixel boundaries for a $250$ ns dwell time, and the stepped gray line indicates the normalised pixel intensity for each pixel with this dwell time. The (normalised) resulting image pixels are shown in the diagram below, for dwell times $\delta_{t}=250,100,50$ ns respectively.
Using a probe-corrected JEOL Grand ARM operated at 300kV, single crystals of [100] oriented SrTiO3 were imaged with a 5.2 pA beam current. Several fast dwell times in the range 50 ns to 1 µs were collected with both with analog and digital signal recording. The microscope used is equipped with fast low-inductance scan coils allowing dwell times as short as $50$ ns to be reached [ishikawa_high_2020]. For the DPC imaging we used a segmented annular all-field (SAAF) detector, which has 16 segments spread over four concentric rings with four segments each [shibata_new_2010]. Performing the in-hardware digitisation has to be done on a per-segment basis and thus one channel per segment must be provided. For the experiments presented here, a four channel Pulse digitiser produced by turboTEM was used and the second ring from the center was chosen for the DPC imaging (see Supplementary Information). Digitisation of 8, 12, or 16 channels is possible by using additional Pulse digitisers in parallel so long as the scan unit can record their outputs (in this case a Gatan Digiscan 3 accepting four simultaneous channels was used). The outer collection angle of the detector annulus used was $32$ mrad and the inner collection angle $16$ mrad, for an $\alpha=30$ mrad beam. A Pulse digitiser is easily retrofittable to any STEM system with BNC cables, and the main demand is that the scan generator accepts the number of channels desirable – preferably directly as digital inputs. Both analog and digitised signals may be recorded simultaneous by forking the signal physically with a T-piece, provided that the number of input channels is sufficient. In order to facilitate multiple, parallel signal digitisation and multi-frame imaging, both analog and digital signal inputs are beneficial as well as a frame clock to aid multi-framing. A screenshot of the software for controlling the Pulse digitiser can be seen in the Supplementary Information.
@@
The in-hardware digitisation calculates the gradient of the PMT signal and thresholds this to count the electron events. This results in a digitised signal which has a true zero noise floor, and where each electron is detected in the correct pixel only and all have the same intensity of unity. Digitising the signal from scintillator detectors has previously been demonstrated to significantly improve signal to noise ratio in HAADF images [mullarkey_development_2021, peters_electron_2023]. Since the digitised pulse appears in the correct temporal position in the PMT stream it also appears in the correct spatial position in the image, without affecting subsequent pixels. As a result, all streaking artefacts are removed. Incidentally, there is some unavoidable few nanoseconds progressing time required by the digitisation hardware, but this manifests as a uniform apparent translation of the sample to the right by some constant fraction of a pixel width. Since scintillator detector surfaces are typically inhomogenous, the electrons may be detected at different intensity depending on where on the detector surface they hit [macarthur_how_2014, sang_characterizing_2016]. Consequently, digitisation of the analog signal ensures a more homogeneous detector response, and yields an image that is quantitative, in number of electrons [macarthur_how_2014, mullarkey_how_2023]. This ensures that detector inhomogeneity does not affect the COM calculations, and that denoising algorithms developed for quantised signals [kusumi_new_2024] may be used on the individual segment images before computing the COMs (although no denoising was used in this study). As a final advantage, having digitised signals makes quantitative comparison with simulations easier.
@@
Figure 6: PSNR in iCOM images for various summing of multi-frame stacks. Time per image is then given as the combined time it took to acquire all the frames, e.g. for 20 summed frames the time per image is $20\cdot T_{\mathrm{Frame}}$. The images on the right are the result of summing 10, 20 and 32 frames respectively for the three different dwell times $\delta_{t}=0.2,0.1,0.05$ µs. The top images are raw summed images and the bottom images are high-pass filtered.
@@
In the graph in Figure 6a we can see how the PSNR increases with increasing ”time per image”, which indicates that the more frames are summed the higher the PSNR in the final analysable image. It is also evident that for a given ”time per image”, the PSNR is generally higher if the frames are captured with a slower dwell time. While this might seem somewhat discouraging for adopting faster scan speeds, a major factor in this is the fly-back time. The faster the scan, the more frames are required to get enough signal for an analysable image, and the more times the beam must fly back. As a result, the time lost to line and frame fly-back is more significant for the faster scan and the temporal resolution achievable from the multi-frame stack may be worsened. In summary, it is vital in these experiments to balance the need for low beam current for viable digitisation, fast scanning for dose fractionation and temporal resolution, and fly-back time considerations. As a final note, with faster scintillators the practicable beam current is increased and with lower-hysteresis scan coils or fly-back compensation [mullarkey_using_2022] the temporal resolution may be increased. Both of which, although beyond the scope of this work, may in the future increase the fidelity and practicability of digitised DPC and improve the temporal resolution for in-situ phase characterisation.